OpenStudy (anonymous):

calculus help. find the derivative of the function. g(x)=(x+5/x^2+6)^6

OpenStudy (anonymous):

which way

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[(\frac{x+5}{x^2+6})^6\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

use the chain rule

OpenStudy (anonymous):

okay the problem is that I did it i mean solves it using the chain rule and quotient but my teacher gave me 6(6-10x-x^2)(5+x)^5/(6+x^2)^7 as the answer....:[

OpenStudy (anonymous):

alright so first you do the derivative of the outside which is \[u^6\] so you get \[6u^5\]

OpenStudy (anonymous):

then the derivative of the inside is the quotient or product rule

OpenStudy (anonymous):

which one would you like to do?

OpenStudy (anonymous):

either one is fine

OpenStudy (anonymous):

alright so you have \[\frac{h(x)g'(x)-g(x)h'(x)}{(h'(x6)^2)}\]

OpenStudy (anonymous):

without the prime at the bottom

OpenStudy (anonymous):



OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so h(x)= \[x^2+6\] and g(x)= \[x+5\] \[h(x)g'(x)=g(x)h'(x)= (x^2+6)(1)-(x+5)(2x)\]

OpenStudy (anonymous):

yup :]

OpenStudy (anonymous):

so now if you multiply everything out you're left with the top like this \[x^2+6-2x^2-10x\]

OpenStudy (anonymous):

yes :]

OpenStudy (anonymous):

combine like terms you get \[ \[-x^2-10x+6\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so now the inside is \[\frac{-x^2-10x+6}{(x^2+6)^2}\]

OpenStudy (anonymous):

okay got it

OpenStudy (anonymous):

i guess my question is what did u do w. the ^6. like after u got the derivative of the inside did u just use the chain rule for it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cause now all together you have \[(\frac{x+5}{x^2+6})^5*\frac{-x^2-10x+6}{(x^2+6)^2}\]

OpenStudy (anonymous):

then you can take the expnent 5 and distribute it to both the numerator and denominator

OpenStudy (anonymous):

technically if you didn't want to do the chain urle you can start by doing this \[\frac{(x+5)^6}{(x^2+6)^6}\]

OpenStudy (anonymous):

and then use the quotient

OpenStudy (anonymous):

but then you'd have to do two chain rules for the quotient

OpenStudy (anonymous):

wouldn't u only multiply by g(x) if we were doing implicit differentiation?

OpenStudy (anonymous):

?

OpenStudy (radar):

Where are you doing "implicit" differentiation?

OpenStudy (anonymous):

never mind :l i am mistaken